EXPLORING POSSIBLE HIGH FUSION POWER REGIMES WITH THE IFS-PPPL MODEL

G. Hammett (PPPL), W. Dorland (U. Md), M. A. Beer (PPPL), M. Kotschenreuther (U. Texas)

Workshop on Burning Plasma Sciences, December, 2000

Extension of paper presented at 1999 Snowmass workshop, http://www.ap.columbia.edu/SMproceedings see also PPPL-3360 (1999):

http://www.pppl.gov/pub report/1999/PPPL-3360-abs.html